OpenStudy (anonymous):

solve for y 3xsquared(3-5y)+sinx=3xy=8

OpenStudy (anonymous):

+ 8

OpenStudy (anonymous):

Clarify.

OpenStudy (anonymous):

solve for y 3xsquared(3-5y)+sinx=3xy+8

OpenStudy (anonymous):

9x^2 -15y + sinx = 3xy +8 9x^2 +sinx -8 = 3xy + 15y = 3y(x+5) y = 9x^2 -8 + sin x /3(x+5)